I need help writing an equation of the line passing through (3,-9) and parallel to the line of 2x+3y=-12. The equation has to be in both point slope and y intercept form
Answer by htmentor(1343)

Parallel lines have the same slope. Putting the equation for the line 2x + 3y = -12 in standard form gives y = (-2/3)x - 4
So the point-slope form of the line we want is y + 9 = (-2/3)(x - 3)
And in intercept form the equation is y = (-2/3)x + 2 - 9 -> y = (-2/3)x - 7